# Diffusion Meets DAgger: Supercharging Eye-in-hand Imitation Learning

### v1

@@ III-C2 Labeling Generated Images
Figure 5: Training Examples from the Diffusion Model: We visualize generated examples, $\tilde{I}$, used to train our policies along with the computed action label (the arrow is a projection of the 3D action into the 2D image plane: the arrow pointing up means move the gripper forward, pointing to the right means move it right). We show augmentations generated by our diffusion model for at different stages of the task. The first row shows augmentations from images where the grabber approaches the apple, while the second row shows augmentations from images where the grabber moves the apple towards the goal location.
Figure 6: The overshooting problem: Under certain conditions, the inferred action for $\tilde{I}_{t}$ may direct the agent away from task success. We refer to this as the overshooting problem. At time step $t+1$, the view $I_{t+1}$ in (c) has moved to the right from $I_{t}$ in (a). However, the synthesized sample $\tilde{I}$ (b) has moved even further to the right than $I_{t+1}$. In (b), Cyan arrow represents action label for $\tilde{I}_{t}$ computed using $I_{t+1}$ as the target; green arrow represents action label computed using $I_{t+3}$ as the target. Since $\tilde{I}_{t}$ has “overshot” $I_{t+1}$, an action taken with $I_{t+1}$ as the next intended target moves backward, away from the goal. Computing the action with respect to a farther image, say $I_{t+3}$ in (d), does not have this issue.
For each image $\tilde{I}_{t}=f(I_{t},\Delta p)$ generated from an original image $I_{t}$, we use $I_{t+k}$ as the target for generating the action label. The action label for $\tilde{I}_{t}$ is simply the action that would convey the agent from the pose depicted in $\tilde{I}_{t}$ to the pose in $I_{t+k}$.
From COLMAP, we obtain camera pose information ${}_{t}T_{w}$ for time step $t$ and ${}_{t+k}T_{w}$ for time step $t+k$. The synthesized view $\tilde{I}_{t}$ can be treated as an image taken by a virtual camera whose pose is represented as ${}_{t}T_{\tilde{t}}$. The diffusion model synthesizes perturbed view conditioned on $I_{t}$ and ${}_{t}T_{\tilde{t}}\;$, or $\tilde{I}_{t}=f(I_{t},{}_{t}T_{\tilde{t}})$ (Section III-B). The action label $\tilde{a}_{t}$ for $\tilde{I}_{t}$ is then computed as $\tilde{a}_{t}={}_{\tilde{t}}T_{t+k}\;=\;{}_{\tilde{t}}T_{t}\,{}_{t}T_{w}\,({}_{t+k}T_{w})^{-1}$. Examples of $(\tilde{I}_{t},\tilde{a}_{t})$ are shown in Figure 5.
One might wonder why we don’t just use $k=1$? As the scale of each trajectory is different, the diffusion model learns the distribution of scales over $\mathbb{D}$. Depending on the scale of the COLMAP reconstruction, this causes the generated samples to move more than 1 unit for some trajectories. When using $k=1$, some generated samples move past the target image $I_{t+1}$ (as described in Figure 6). This causes conflicting supervision, as the computed action does not make progress toward completing the task. Using a large $k$ guards against this, leading to stronger performance downstream, as our experiments verify (Section IV-C2).
Figure 7: Robotic Platform for Online Testing: We conduct our experiments on a Franka Research 3 robot with a wrist-mounted GoPro Hero 9. In the depicted configuration, the task is to push the apple onto the red circle. Since VIME’s [77] action space is relative 3D translations, we modified its grabber mount on the robot, allowing the robot to move without reaching joint limits while preserving the original rotation of the end-effector.
@@ IV Experiments
IV-A Experiment Setup
IV-A1 Task, Observation Space, Action Space
We use the non-prehensile pushing task from VIME [77] as a benchmark task. The task is to push the target object to a target location marked by a red circle and requires both reaching and pushing the object. Following Young et al. [77], we use a grabber as the robot end-effector as shown in Figure 7. This allows for easy collection of expert demonstrations using a GoPro mounted on a Grabber. Due to the difference in joint limits between the XArm (used by VIME) and the Franka Research 3 robot, we redesign the attachment between the grabber and the robot flange. In this way, the robot can execute planar push in the typical top-down configuration, extending the workspace to a wider range of the table. The comparison is shown in Figure 7. Observations come from the eye-in-hand GoPro camera. We use the action space used in VIME’s [77] publicly available code: relative 3D translations of the end-effector.
IV-A2 Task and Play Data
We collect task and play data trajectories using a GoPro mounted on a Grabber. For task data, an expert pushes the object to a target location. For play data, we move the grabber around the apple and push it in many different directions. Through this procedure, we only get image sequences. We use structure from motion [21, 68] to extract camera poses ${}_{t}T_{w}$ for images in the collected sequences. Action $a_{t}$ is computed as the relative translation between $I_{t+1}$ and $I_{t}$ in $I_{t}$’s camera frame. Since COLMAP only gives reconstructions upto an unknown scale factor, following [77] we normalize the action to unit length and interpret it as just a direction. Since the final robot action space is 3D translation, we avoid rotating the grabber during both type of data collection; however, slight rotation of the grabber still exists in expert demonstrations.
IV-A3 Policy Architecture and Execution
@@
IV-A4 Comparisons
We use behavior cloning on the expert data as done in past work [77], as the baseline. We refer to this as BC. We also compare to SPARTN [85] and evaluate the various design choices. Since there is no publicly available code from SPARTN, we replicate their procedure using NerfStudio [67].
IV-B Visual Comparison of Generated Images
We start by assessing the visual quality of samples generated by the diffusion model in DMD. Figure 4 and Figure 5 shows sample images generated by our diffusion model. The model is able to faithfully generate images where the camera is moved left / right and front / back with minimal artifacts.
We also compare the visual quality of our generated images to those generated using the NeRF based approach from SPARTN [85]. The moving gripper in the scene breaks the static scene assumption made by NeRF. Thus, [85] mask out the gripper before training the NeRF. While this works for the pre-grasp trajectory that Zhou et al. [85] seek to imitate, this strategy fails when the gripper actually manipulates the scene, as in our task. Thus, we investigate different masking schemes: a) no masking, b) masking a larger region around the gripper, c) masking just the gripper; and show visualizations in Figure 8. No masking leads to the worst results as expected. Masking just the grabber fixes the grabber, but causes the object being manipulated to blur out. Masking a larger region around the grabber fixes the blurring but also eliminates the object being manipulated. In contrast, images synthesized using our diffusion model move the camera as desired without creating any artifacts.
Figure 8: Diffusion vs NeRF Here we visualize perturbed samples generated using DMD and NeRF baselines with different masking strategies. The top row shows images generated corresponding to a forward movement relative to $I_{t}$, and the bottom row for a backward movement. When using a NeRF with no masking (a), the reconstructions of the gripper and apple are degenerate, as they violate the static scene assumption in NeRFs. With a box-mask (b), the apple is occluded by the mask, and consequently, missing from the reconstruction. Even with a mask around the end-effector (Grabber-Mask), there are major artifacts in reconstructing the apple (c). With ours (Diffusion Model), the model faithfully reconstructs the grabber and apple (d).
IV-C Offline Validation
We collected 74 expert demonstrations in total for the apple pushing task. We use 24, 23, and 27 trajectories for train, validation, and test, respectively. We use vanilla behavior cloning on the expert data as the baseline as done in past work [77]. Our offline validation involves evaluating the policy’s predictions on the test set. While the test set samples doesn’t contain examples of going off the distribution, it does measure generalization to some extent by evaluating the policy on states outside of the training set. For the error metric, we use the median angle between the predicted and the ground truth translation.
We conduct offline validation to compare with other data augmentation schemes and select the value of $k$ to mitigate the over-shooting problem. Online experiments in Section IV-D test the success rate of the policy.
@@ IV-C Offline Validation
IV-C1 Comparison with other data augmentation schemes
We want to understand the effectiveness of augmenting with diffusion-generated images compared to augmenting with standard techniques, such as color jitter and horizontal flip. Table I shows that augmentation, in general, improves performance, and augmenting with out-of-distribution images leads to larger improvement compared to standard techniques. Comparing the two columns, we see that adding diffusion-generated images improve performance in all settings.
IV-C2 Overshooting problem, what frame to use?
@@
Figure 9: Effect of using different future frames for labeling augmenting images on translation prediction: to overcome the overshooting problem described in Figure 6, we experiment with different future frame $I_{t+k}$ for labeling the diffusion model generated images. Error decreases as $k$ increases and plateaus around $k=3$. Therefore, we use frame $I_{t+3}$ for labeling $\tilde{I_{t}}$
@@
In order to mitigate the overshooting problem, we use image $I_{t+k},k>1$ to label the augmenting image instead of $I_{t+1}$. We evaluate policies trained with labels produced from different $k$’s, and show the results in Figure 9. We find that labeling with $I_{t+1}$’s pose works worse than labeling with pose for $I_{t+3}$ and beyond. As the curve plateaus beyond $k=3$ and using larger $k$ leads to the exclusion of the last $k-1$ frames in each trajectory, we use $k=3$ for our online experiments.
Table II: Online Robot Experiment Results. We conduct 5 experiments: 10 trials per method and randomize the apple start locations between trials.
IV-D Online Validation
@@
Figure 10: Comparison of different methods on staying on course. We show the trajectories executed by BC, SPARTN [85], and DMD over several steps towards the target. Unlike BC and SPARTN, which gradually deviate from the intended path, DMD maintain its course towards the right and reach the target successfully. See project website for more execution videos.
@@
Our online experiments measure the success rate at the non-prehensile pushing task. We conduct 10 trials per method and randomize the apple start location between trials. To prevent human biases, we choose the apple start location before sampling the method to run next. Execution trajectories can be found on the project website.
IV-D1 DMD vs. Behavior Cloning
@@
We take the best model for BC and DMD from Table I and evaluate them on the robot. We find the vanilla BC model to get a 30% success rate while DMD achieves 100% (Table IV-C2 a). This significant difference shows that minor error at each step can accumulate and result in task failure.
IV-D2 DMD with Few Demonstrations
@@
Motivated by the 100% performance of DMD with 24 demonstrations, we ask how low can the number of demonstrations be? We consider 2 low data settings with 8 and 16 demonstrations. For these settings, the diffusion model is trained with only play data. Success rate for these two models on the robot is shown in Table IV-C2 c. Once again DMD outperforms BC and achieves 80% success even when trained with only 8 demonstrations.
IV-D3 DMD vs. SPARTN [85]
@@
Following up on the visual comparison of the quality of the augmenting samples from Section IV-B, we conduct a head-to-head comparison against SPARTN on the real robot. SPARTN achieves a success rate of 50% vs. DMD succeeds 100% of the time (Table IV-C2 b). In Figure 10, we show that while SPARTN diverges from the expert course and pushes the apple off the table, DMD brings the apple to the target sucessfully.
IV-D4 Utility of Play Data for Training Diffusion Model
@@
We also experimentally validate what the choice of the dataset necessary for training an effective diffusion model for view synthesis. We repeat the experiment with 24 demos but vary the diffusion model that generated the augmenting samples. Using the diffusion model trained on just task data lead to an 80% success rate vs. 100% success rate with the diffusion model trained on a combination of task and play data (Table IV-C2 d). Thus, while just training the diffusion model on task data is effective (a nice result in practice), performance can be boosted further by training the diffusion model on play data.
IV-D5 Scaling to Other Objects
We conduct an experiment on another object, a cup. The BC baseline achieves a success rate of 80% vs DMD achieves a success rate of 90% (Table IV-C2 e).

### v2

@@ III-C2 Labeling Generated Images
For each image $\tilde{I}_{t}=f(I_{t},\Delta p)$ generated from an original image $I_{t}$, we use $I_{t+k}$ as the target for generating the action label. The action label for $\tilde{I}_{t}$ is simply the action that conveys the agent from the pose depicted in $\tilde{I}_{t}$ to the pose in $I_{t+k}$.
Using SfM, we obtain camera pose information ${}_{t}T_{w}$ for time step $t$ and ${}_{t+k}T_{w}$ for time step $t+k$. The synthesized view $\tilde{I}_{t}$ can be treated as an image taken by a virtual camera whose pose is represented as ${}_{t}T_{\tilde{t}}$. The diffusion model synthesizes perturbed view conditioned on $I_{t}$ and ${}_{t}T_{\tilde{t}}\;$, or $\tilde{I}_{t}=f(I_{t},{}_{t}T_{\tilde{t}})$ (Section III-B). The action label $\tilde{a}_{t}$ for $\tilde{I}_{t}$ is then computed as $\tilde{a}_{t}={}_{\tilde{t}}T_{t+k}\;=\;{}_{\tilde{t}}T_{t}\,{}_{t}T_{w}\,({}_{t+k}T_{w})^{-1}$. Examples of $(\tilde{I}_{t},\tilde{a}_{t})$ are shown in Figure 4.
One might wonder why we don’t just use $k=1$? When sampling $\Delta p$, we assume that the perturbed observation $\tilde{I}_{t}$ is within a local region around $I_{t}$. However, some generated samples can move past the target image $I_{t+1}$ (described in Figure 5). This causes conflicting supervision, as the computed action does not make progress toward completing the task. Using a large $k$ guards against this, leading to stronger performance downstream, as our experiments verify (Section IV-A3).
Figure 6: DMD Robotic Experiments: There are four tasks in total: pushing apple to target location (Section IV-A), stacking five different cups on box (Section IV-B), pouring coffee beans into a cup (Section IV-C), and hanging shirt on a rack (Section IV-D). We conduct our experiments on a Franka Research 3 robot with a wrist-mounted GoPro Hero 9. We modified VIME’s [78] grabber mount for Franka, allowing the robot to reach end-effector poses without reaching joint limits.
@@ IV Experiments
We test DMD across four tasks: non-prehensile pushing, stacking, pouring, and hanging a shirt. The task setups are shown in Figure 6. Together these tasks test DMD in different settings: 3DoF action space (pushing, stacking), 6DoF action spaces (pouring, hanging a shirt), generalization to new objects (stacking), precision in reaching goal locations (pouring, hanging a shirt).
On the pushing task, we present visual comparisons to NeRF-based synthesis approach SPARTN [86] in Section IV-A2 and in-depth quantitative analysis (ablation of design choices, offline evaluations) in Section IV-A3. We then show a variety of online comparisons for the pushing task in Section IV-A4. Moreover, we compare DMD to behavior cloning for stacking (Section IV-B), pouring (Section IV-C), and hanging a shirt (Section IV-D). Finally, we test whether DMD improves generalization to novel objects and environment when provided with a diverse task dataset, as described in Section IV-E. We adopt randomized A/B testing when making comparisons between different methods to minimize the effects of unknown environmental factors.
Figure 7: Visualization of Perturbed Images Generated by Diffusion Model for Augmenting Policy Learning: In each row, we show $\tilde{I_{t}}$ generated from $I_{t}$ through different camera translations for the tasks of pushing, stacking, pouring, and hanging a shirt.
IV-A Non-prehensile Pushing
IV-A1 Experiment Setup
Task, Observation Space, Action Space
We use the non-prehensile pushing task from VIME [78], which involves reaching a target object and pushing it to the target location marked by a red circle. Following Young et al. [78], we use a grabber as the robot end-effector as shown in Figure 6. This allows for easy collection of expert demonstrations using a GoPro mounted on a grabber. Due to the difference in joint limits between the XArm (used by VIME) and the Franka Research 3 robot, we redesign the attachment between the grabber and the robot flange (Figure 6(a)). In this way, the robot can execute planar push in the typical top-down configuration, extending the workspace to a wider range of the table. Observations come from the eye-in-hand GoPro camera. For pushing, we use the action space used in VIME’s [78] publicly available code: relative 3D translations of the end-effector.
Task and Play Data
We collect task and play data using a GoPro mounted on a grabber. The task data include 74 demonstration, and the play data include 36 trajectories. For task data, an expert pushes the object to a target location. For play data, we move the grabber around the apple and push it in different directions. From the GoPro videos, we obtain image sequences and use structure from motion [23, 69] to extract camera poses ${}_{t}T_{w}$. Action $a_{t}$ is computed by computing the relative translation between $I_{t+1}$ and $I_{t}$ in $I_{t}$’s camera frame. Since COLMAP only gives reconstructions up to an unknown scale factor, we normalize the action to unit length and interpret it as just a direction.
Policy Architecture and Execution
@@ IV-A1 Experiment Setup
Baselines
We use vanilla behavior cloning on the expert data as the baseline as done in past work [78], we refer to this as BC. We also compare to SPARTN [86] and evaluate the various design choices. Since there is no publicly available code from SPARTN, we replicate their procedure using NerfStudio [68].
IV-A2 Visual Comparison of Generated Images
Figure 7 shows the high visual quality of samples generated by the diffusion model. The model is able to faithfully synthesize images where the camera is moved left / right, up / down, and front / back with minimal artifacts.
As shown in Figure 9, we also compare the quality of diffusion-generated images to those generated using the NeRF based approach from SPARTN [86]. The moving gripper breaks the static scene assumption made by NeRF. Thus, [86] masks out the gripper before training the NeRF. While this strategy works for the pre-grasp trajectory that Zhou et al. [86] seek to imitate, it fails when the gripper manipulates the scene, as in our tasks. Thus, we investigate different masking schemes: a) no masking, b) masking just the gripper, c) masking a larger region around the gripper; and show visualizations in Figure 9.
No masking leads to the worst results as expected. Masking just the grabber fixes the grabber, but causes the object being manipulated to blur out. Masking a larger region around the grabber fixes the blurring but also eliminates the object being manipulated. In contrast, images synthesized using our diffusion model move the camera as desired without creating any artifacts.
@@ IV-A2 Visual Comparison of Generated Images
Figure 8: Effect of Using Different Future Frames for Labeling Augmenting Images: We experiment with using different future frame $I_{t+k}$ for labeling the diffusion-generated images. Error decreases as $k$ increases and plateaus around $k=3$. Therefore, we use frame $I_{t+3}$ for labeling $\tilde{I_{t}}$
Figure 9: Diffusion vs NeRF We visualize perturbed samples generated using DMD and NeRF with different masking strategies. The top row shows images generated for a forward movement relative to $I_{t}$; the bottom row shows images for a backward movement (a) With no masking, the NeRF reconstructions of the gripper and apple are degenerate, as they violate the static scene assumption in NeRFs. (b) With a box-mask, the apple is occluded by the mask, and consequently, missing from the reconstruction. (c) Even with a mask around the end-effector (Grabber-Mask), there are major artifacts in reconstructing the apple. (d) With ours (Diffusion Model), the model faithfully reconstructs the grabber and apple.
IV-A3 Offline Validation
Out of the 74 expert demonstrations, we use 24, 23, and 27 trajectories for train, validation, and test, respectively. As done in past work [78], the baseline is vanilla behavior cloning on the expert data. Our offline validation evaluates the policy’s predictions on the test set using the median angle between the predicted and ground truth translations as the error metric.
Comparison with other data augmentation schemes.
The goal of this experiment is to compare the effectiveness of augmenting with diffusion-generated images to standard techniques, such as color jitter and horizontal flip. Table I shows that augmenting with out-of-distribution images leads to larger improvement compared to standard techniques. Comparing the two columns, we see that adding diffusion-generated images improve performance in all settings.
Table II: Pushing Online Evaluation Results. (a) DMD outperforms BC across all settings. DMD achieves a 100% success rate when pushing an apple, greatly exceeding BC’s 30%. It also maintains an 80% success rate with only 8 demonstrations, whereas BC drops to 20%. (b) As shown in Figure 9, our diffusion model synthesizes higher quality images than NeRFs, especially when scenes undergo deformations. This advantage results in higher task performance: DMD achieves a 100% success rate, while SPARTN [86] achieves only 50%. (c) Training the diffusion model with additional play data boosts the task success rate to 100%, compared to 80% when using the model trained only on task data.
(a) DMD vs. BC
[…]
(b) DMD vs. SPARTN [86]
[…]
(c) Utility of Play Data
Method Apple DMD Task Only 80% DMD Task & Play 100%
Overshooting problem, what frame to use?
@@ IV-A3 Offline Validation
In order to mitigate the overshooting problem, we use image $I_{t+k},k>1$ to label the augmenting image instead of $I_{t+1}$. Figure 8 shows the effect of this parameter $k$ on policy performance. We find that labeling with $I_{t+1}$’s pose works worse than labeling with $I_{t+3}$’s pose and beyond. As the curve plateaus for $k>3$ and using larger $k$ leads to the exclusion of the last $k-1$ frames in each trajectory, we use $k=3$ for our online experiments.
IV-A4 Online Validation
@@
Figure 10: Comparison of BC, SPARTN, and DMD for Staying on Course. We show the trajectories executed by BC, SPARTN [86], and DMD over several steps towards the target. Unlike BC and SPARTN, which gradually deviate from the intended path, DMD maintain its course towards the right and reach the target successfully. See project website for more execution videos.
@@
We conduct 10 trials per method and randomize the apple start location between trials. To prevent human biases, we choose the apple start location before sampling the method to run next. Execution trajectories can be found on the project website.
DMD vs. Behavior Cloning
@@
We take the best model for BC and DMD from Table I and evaluate them on the robot. We find the vanilla BC model to get a 30% success rate while DMD achieves 100% (Table II(a)). This significant difference shows that minor error at each step can accumulate and result in task failure.
DMD with Few Demonstrations
@@
Motivated by the 100% performance of DMD with 24 demonstrations, we ask how low can the number of demonstrations be? We consider 2 low data settings with 8 and 16 demonstrations. For these settings, the diffusion model is trained with only play data. Success rate for these two models on the robot is shown in Table II(a). Once again DMD outperforms BC and achieves 80% success even when trained with only 8 demonstrations.
DMD vs. SPARTN [86]
@@
Following up on the visual comparison of augmenting samples from Section IV-A2, we conduct a head-to-head comparison against SPARTN on the real robot. SPARTN achieves a success rate of 50% vs. DMD succeeds 100% of the time (Table II(b)). In Figure 10, we show that while SPARTN diverges from the expert course and pushes the apple off the table, DMD brings the apple to the target successfully.
Utility of Play Data for Training Diffusion Model
@@
We also experimentally validate what the choice of the dataset necessary for training the view-synthesis diffusion model. We repeat the experiment with 24 demos but vary the diffusion model that generated the augmenting samples. Using the diffusion model trained on just task data lead to an 80% success rate vs. 100% success rate with the diffusion model trained on a combination of task and play data (Table II(c)). Thus, while just training the diffusion model on task data is effective (a nice result in practice), performance can be boosted further by training the diffusion model on play data.
Scaling to Other Objects
We test DMD on another object, a cup. The diffusion model is trained with only cup play data. Task data include 16 demonstrations. BC is trained only with task data, and DMD is trained additionally with synthesized images. Similar to apple, we repeat each method 10 times and randomize the cup start location. We choose the start location before sampling the method to execute. The BC baseline achieves a success rate of 80% vs DMD achieves a success rate of 90% (Table II(a)).
IV-B Stacking
IV-B1 Experiment Setup
Stacking involves reaching a cup, grasping it, placing it on a black box, and opening the grabber. This task is similar to the stacking task in VIME [78] and requires precision in the height direction to move the cup on the box. The action space is the same as in VIME’s publicly available code: relative 3D translations of the end-effector.
We use a grabber and collect 50 demonstrations across three training cups (red dotted box in Figure 6(b)). We obtain cameras poses using ORB-SLAM3 [8, 15], which recovers the correct metric scale by using the IMU data from GoPro.
We train the diffusion model on the 50 expert trajectories and don’t use any play data. Unlike pushing, we do not need to scale by the largest displacement between adjacent frames because of metric scale from IMU. To sample out-of-distribution images, we sample a small translation (with length between $2cm$ and $4cm$) in a random direction and a small rotation obtained by sampling yaw, pitch, and roll uniformly from the range $[-10^{\circ},10^{\circ}]$, $[-10^{\circ},10^{\circ}]$, and $[-10^{\circ},15^{\circ}]$, respectively.
The policy architecture is the same as that of pushing and is trained with L1 loss to the normalized ground-truth translation. We annotate frames at which the grabber open/close and train a gripper policy that is shared across all the methods.
IV-B2 Online Validation
We evaluate on the three training cups plus two holdout cups: a larger red cup and a Sprite can (yellow dotted box in Figure 6(b)). The starting locations of the objects are randomized and do not necessarily match the training data.
Table III: Stacking Online Evaluation Results. In the task of stacking a cup on a box, DMD achieves over 90% success rate for all training cups and over 80% for the two cups unseen cups and outperforms BC. The diffusion model is only trained on task data/expert demonstrations. See all execution videos on project website.
Seen Cups Unseen Cups
Method Blue Red-1 White Red-2 Sprite
BC 10% 10% 90% 40% 50%
DMD (Task Only) 100% 90% 100% 80% 90%
As shown in Table III, DMD performs better than BC across both seen and unseen instances. BC often fails to lift the tall cups above the box and pushes the box forward continuously. See videos on project website for failure modes. This behavior might explain the higher success rate of the white cup, which is shorter and thus easier to lift above the box. By training on synthesized data, DMD learns to lower the grabber if approaching the cups too high, and it grasps closer to the cup bottom so that moving above the box becomes easier. It succeeds more than BC across all cups.
IV-C Pouring
IV-C1 Experiment Setup
As shown in Figure 6(c), this task requires the robot to reach a blue cup filled with 57 grams of coffee beans, grasp it, and pour all the beans into a red cup. Success is determined by the weight of the coffee beans being the same at the end of each trial. Action space is full 6-DoF.
Task data contain 49 demonstrations, and play data contain 16 trajectories. Diffusion model is trained on both types of data. Except for an additional MLP-head to predict rotation, the policy architecture is the same as that of stacking; the model is trained with L1 loss to the ground-truth translation and rotation.
During execution, the robot executes the translation and rotation predicted by the model. Locations of the two cups are randomized, but the blue cup is always on the same side of the red cup.
Table IV: Pouring and Hanging-Shirt Online Evaluation Results. DMD outperforms BC across both tasks. See all execution videos on project website. : the two BC numbers are different because they are from two different pairwise randomized A/B tests: 1) BC vs. DMD (Task & Play) and 2) BC vs. DMD (Task Only), and experimental conditions (e.g. lighting) may have been different between when these two tests were conducted.
Pouring Hanging
Method Task & Play Task Only Task & Play
BC $0\%^{\ast}$ $30\%^{\ast}$ 20%
DMD 80% 70% 90%
IV-C2 Online Validation
DMD succeeds 80% of the time while BC fails completely. A common failure case for BC is that as the robot rotates the cup with coffee beans, it does not move the cup closer to the receiving cup; the blue cup then drifts further and further from the one on the table. In contrast, even in the worst trial, DMD can pour 53 grams (i.e. more than 90%) of the coffee beans. See execution videos on project website for clear differences between BC and DMD.
We also validate the use of play data for learning an effective diffusion model. We finetune the diffusion model with only task data and conduct the same experiment as the one in the last paragraph. As shown in Table IV, without training the diffusion model with play data, the downstream policy maintains a 70% success rate, higher than BC’s 30%. This shows that, training on task data alone also enables the diffusion model to synthesize good out-of-distribution images that improve task performance. Note that the two BC numbers are different because they are from two different pairwise randomized A/B tests: 1) BC vs. DMD (Task & Play) and 2) BC vs. DMD (Task Only), and experimental conditions (e.g. lighting) may have been different between when these tests were conducted.
IV-D Hanging a Shirt
IV-D1 Experiment Setup
To hang the shirt, the robot reaches the hanger and slides the shirt across the table to the edge; it then grasps the part of the hanger sticking out of the table and hangs the shirt on the clothing rack. The setup is shown in Figure 6(d). Action space is full 6-DoF.
Task data contains 55 demonstrations, and play data contains 24 trajectories. The diffusion model is trained on both types of data. We divide the hanging task into two sub-tasks: grabbing the shirt and putting it on the rack. A separate policy is learned for each task. We use same policy architecture and loss function as those used in pouring.
During execution, the robot starts with the grabbing policy. When the network predicts to close the grabber, the grabbing policy terminates, and the robot switches to the putting-on-rack policy. The entire process terminates when the top of the hanger makes contact with the rack pole, signaled by when the end-effector senses a force greater than $4N$. We add some clay padding to the grabber tips because the hanger is thin, and the motor is not strong enough to close the grabber completely.
IV-D2 Online Validation
As shown in Table IV, DMD achieves a 90% success rate while BC only succeed in 20% cases. BC often fails because the robot raises the shirt too high, and the shoulder part of the hanger hits the pole before the hook part reaches the pole. On the other hand, DMD learns to bring the shirt down onto the rack.
IV-E In-the-Wild Cup Arrangement
Figure 11: In-the-Wild Cup Arrangement Setup: We evaluate DMD on the in-the-wild cup arrangement task from UMI [15]. The training set for this task contains 1447 demonstrations across 30 locations and 18 cups. Because our pre-trained diffusion model operates on non-fish eye images, we work with center-cropped and rectified images from the Fisheye lens. Baseline and baseline+DMD operate on these center-cropped rectified images. Evaluation is done on 5 novel cups in a novel environment (our lab).
Table V: In-the-Wild Cup Arrangement Online Evaluation Results. DMD generalizes better to novel cups in novel environment than vanilla Diffusion Policy [14], achieving a 64% overall success rate compared to Diffusion Policy’s 16%. It succeeds in same or more trials across all 5 cups. See all execution videos on project website.
Method Navy White Mint Red Blue Overall
Diffusion Policy [14] 1/5 1/5 1/5 0/5 1/5 4/25 (16%)
DMD (Task Only) 3/5 3/5 3/5 3/5 4/5 16/25 (64%)
The goal of this experiment is to understand a) can DMD improve performance in the presence of a large number of diverse demonstrations, b) can DMD improve generalization of policies to novel objects in novel environments, and c) if DMD is compatible with policies trained with diffusion [14].
IV-E1 Experiment Setup
We leverage a diverse in-the-wild dataset from the recent Universal Manipulation Interface (UMI) paper [15]. It introduces a hand-held data collection device and an imitation learning framework utilizing Diffusion Policy [14]. The device captures image observations using a wrist-mounted GoPro camera.
We adopt the same task definition as the in-the-wild generalization experiment in UMI: placing a cup on a saucer with its handle facing the left side of the robot. UMI [15] collected 1447 demonstrations across 30 locations and 18 training cups. We use their publicly available demonstration data and conduct evaluation in our lab (i.e. novel location) with and without DMD.
Our evaluation setup is identical to that used in UMI, with one difference. UMI uses wide field-of-view images obtained by putting an additional Fisheye lens on the GoPro. Our diffusion model denoises in the latent space of a VQ-GAN autoencoder [16] that was pre-trained on the RealEstate10K dataset [87], which contains non fish-eye images. Therefore, our diffusion model is constrained to only work well on non-fish eye images. Thus, we conduct comparisons where both the baseline and baseline+DMD operate on center-cropped rectified images shown in Figure 11. We retrain the baseline diffusion policy from UMI to consume these center-cropped rectified images. Other than this change, the task policy uses UMI’s implementation of Diffusion Policy [14] and denoises 6DoF actions.
We test on 5 held-out cups shown in Figure 11. These cups are not used for expert demonstrations in dataset from [15]. For each cup, we test 5 different start configurations. We follow the experiment protocol outlined in [15]: we use pixel masks to make sure that the starting locations of the cups and saucers are the same across the two methods.
IV-E2 Online Validation
DMD improves upon the vanilla Diffusion Policy [14] across all novel cups, achieving an overall success rate of 64%, compared to Diffusion Policy’s 16% success rate (Table V). This improvement shows that DMD is effective even in the presence of large number (1447) of diverse demonstrations and also leads to improved performance when testing on novel objects in novel environments. Furthermore, DMD-created data also improves policies trained via diffusion. One common failure case for the baseline is that the policy doesn’t reorient the cup correctly, because it fails to make contact with the cup or its handle. DMD, on the other hand, successfully pushes the handle to the left side more often and grasps the cup without tipping it over.